# CoTAL: Human-in-the-Loop Prompt Engineering, Chain-of-Thought Reasoning, and Active Learning for Generalizable Formative Assessment Scoring

### v2

@@ Abstract
Large language models (LLMs) have created new opportunities to assist teachers and support student learning. Methods such as chain-of-thought (CoT) prompting enable LLMs to grade formative assessments in science, providing scores and relevant feedback to students. However, the extent to which these methods generalize across curricula in multiple domains (such as science, computing, and engineering) remains largely untested. In this paper, we introduce Chain-of-Thought Prompting + Active Learning (CoTAL), an LLM-based approach to formative assessment scoring that (1) leverages Evidence-Centered Design (ECD) principles to develop curriculum-aligned formative assessments and rubrics, (2) applies human-in-the-loop prompt engineering to automate response scoring, and (3) incorporates teacher and student feedback to iteratively refine assessment questions, grading rubrics, and LLM prompts for automated grading. Our findings demonstrate that CoTAL improves GPT-4’s scoring performance, achieving gains of up to 24.5% over a non-prompt-engineered baseline. Both teachers and students view CoTAL as effective in scoring and explaining student responses, each providing valuable refinements to enhance grading accuracy and explanation quality.
Index Terms:
@@
Human-in-the-Loop, Formative Assessment, Automated Short Answer Scoring, Automated Grading, Prompt Engineering, LLM, LLMs, K12 STEM.
I Introduction
@@
In K-12 STEM+C (Science, Technology, Engineering, Mathematics, and Computing) classrooms, educators foster engagement by linking scientific principles to real-world phenomena, enabling students to develop exploration, inquiry, and problem-solving skills. This provides students with opportunities to develop a foundational understanding of essential STEM+C concepts and practices [1]. Unlike single-discipline curricula, STEM+C learning requires students to integrate cross-domain concepts synergistically, linking ideas from one domain (e.g., science) to another (e.g., computing). While this approach enhances learning outcomes and promotes a deeper understanding of scientific processes, it also introduces complexities that can hinder learning [2], necessitating additional guidance and support that formative assessments can provide.
@@
Assessments can incorporate open-ended questions that help students identify key learning constructs, apply learned concepts to problem-solving tasks, and develop critical thinking skills [3]. Simultaneously, they provide teachers with deeper insights into students’ STEM+C knowledge and problem-solving abilities [4]. Unlike summative assessments and standardized tests, which are primarily used for evaluation and may not capture the complexities students may face in achieving their learning goals [5], formative assessments are designed to support self-reflection when students have difficulties and facilitate feedback that helps them refine their understanding and improve their performance. At the same time, it allows teachers to monitor students’ learning progress and adjust their instruction to better meet student needs.
@@
Evidence-based approaches for generating formative assessments and evaluation rubrics ensure alignment with curricular goals and designated standards [6]. Evidence-centered design (ECD) enables a more nuanced and flexible approach by adopting cognitive science and instructional design approaches that embed evidentiary reasoning into each stage of assessment and rubric development [5].
@@
However, challenges persist in classroom environments for generating, grading, and providing timely feedback for formative assessments that align with educational standards. Designing and evaluating free-response formative assessments can place excessive demands on classroom teachers, who may have limited expertise in integrating STEM+C subjects, potentially affecting their ability to create assessments and rubrics that accurately evaluate students’ interdisciplinary knowledge [1]. Students may also lack mature writing skills, necessitating significant time and effort from human graders to infer the true implications of students’ answers [7]. Consequently, research is needed to develop automated scoring systems that efficiently deliver needed feedback to facilitate STEM+C learning based on teacher preferences [8].
@@
Prompt engineering with large language models (LLMs) offers a promising solution, allowing users to adapt language models to downstream tasks by incorporating approaches like in-context learning (ICL) [9], chain-of-thought (CoT) prompting [10], and active learning [11] that obviate the need for traditional training via parameter updates, thus conserving computational resources. ICL enables LLMs to “learn” from labeled few-shot examples in the prompt during inference. CoT extends ICL by augmenting labeled few-shot examples with step-by-step reasoning chains to provide more explicit guidance to the LLM for scoring student answers [10]. Active learning is a human-in-the-loop approach to improving model training, where the human serves as an oracle, selecting additional instances to label for the next training iteration to enhance performance and robustness [11].
Unlike algorithmic prompt engineering, human-in-the-loop approaches [12, 13] integrate an important human perspective into the design and optimization pipeline that enable users to influence LLM outputs rather than relying solely on algorithmic decisions [14, 13]. This creates a human-AI collaboration [15] that enhances LLM alignment with user preferences and ensures the contextual relevance of the generated output by incorporating nuanced domain knowledge that algorithmic methods may overlook. Furthermore, this approach helps mitigate LLM errors and hallucinations by leveraging human expertise to validate and refine generated responses, fostering more accurate, ethical, and trustworthy LLM interactions.
@@
In this paper, we pursue a stakeholder-AI partnership [^1] For the purposes of this paper, “stakeholders” refers to teachers, students, and researchers. to enrich assessment and feedback opportunities in a combined science, engineering, and computing NGSS-aligned curricular unit on Earth sciences [16, 17]. Building on prior work in automated scoring for short-answer science assessments, we introduce a generalizable human-in-the-loop LLM prompt engineering approach, Chain-of-Thought Prompting + Active Learning (CoTAL), to automate scoring and feedback for formative assessments across multiple domains. We define “generalizable” as the ability to accurately score formative short-answer responses that differ by question type (e.g., concept definitions, process descriptions, comparisons, and explanations), rubric structure (e.g., multi-label vs. multi-class), and content domain (science, computing, and engineering).
CoTAL consists of three phases: (1) leveraging ECD principles to design curriculum-aligned formative assessments and rubrics; (2) integrating these with human-in-the-loop prompt engineering via ICL, CoT, and active learning; and (3) refining the assessments, rubrics, and grading prompts through stakeholder feedback. This approach enhances LLM-based formative assessment response scoring and explanation capabilities while remaining grounded in principled assessment design.
Within this framework, we address the following research questions:
@@
Our results demonstrate that CoTAL generalizes effectively, significantly enhancing the performance of base GPT-4. Furthermore, CoTAL’s ability to clarify its scoring decisions builds trust in AI systems among students and teachers. This paper provides a pathway for introducing LLM-based assessment scoring in education and the learning sciences. The CoTAL approach can be generalized to show how LLMs can enhance automated formative assessment scoring and feedback generation while aligning with individual teacher preferences. For students, it provides additional learning opportunities by helping them reflect on their answers and overcome their difficulties.
II Background
@@
LLMs offer new and exciting opportunities for addressing formative assessment grading and feedback generation [21, 22]. Previously, automated assessment scoring methods have incorporated data augmentation [23, 24], next sentence prediction [25], domain adaptation and supervised fine-tuning [26], prototypical neural networks [27], cross-prompt fine-tuning [28], and reinforcement learning [29] to improve grading accuracy. However, these methods have largely focused on more structured grading tasks in mathematics and computer science [30, 31], where open-ended responses are less prevalent than in science domains [12]. These approaches have achieved varying degrees of success but often fail to provide comprehensive insights into their scoring decisions.
Recently, researchers have used prompt engineering techniques to improve LLM performance on formative assessment scoring [12, 32]. However, in most cases, these approaches target a single domain and dataset, and their generalizability remains largely unevaluated. These approaches also fail to consider stakeholder input when developing and refining their methods. For example, Lee et al. (2024) [33] demonstrates the effectiveness of CoT prompting using contextual item stems and rubrics, emphasizing the importance of domain-specific reasoning in improving LLM performance when scoring middle school science assessments. Their prompt engineering procedure, WRVRT (Writing, Reviewing, Validating, Revising, and Testing), leverages CoT prompting and iterative refinement to enhance scoring accuracy and explainability. However, their study focuses solely on the science domain, and it remains unclear how well their approach generalizes to other subjects. Their work also centers on the prompt engineering process, without incorporating ECD principles or integrating feedback from students and teachers to refine their systems.
@@
There exists a notable gap in automated formative assessment scoring approaches that (1) integrate ECD design principles with prompt engineering, (2) explain scoring decisions to teachers and students, (3) are effective across multiple domains, and (4) incorporate stakeholder feedback for refinement. In this paper, we address these gaps by proposing a generalizable, stakeholder-informed framework that combines ECD with human-in-the-loop prompt engineering to support transparent, accurate, and adaptable formative assessment scoring. Our work builds on prior research by explicitly grounding our methodology in assessment design theory, engaging teachers and students to refine our approach and evaluating performance across science, engineering, and computing domains.
III SPICE Curriculum
@@
SPICE is a three-week middle school curriculum unit that challenges students to redesign their schoolyard using surface materials to minimize the amount of water runoff after a rainstorm while adhering to design constraints (such as considering construction cost and accessibility needs). The problem-based learning curriculum comprises five core units: (1) physical experiments; (2) conceptual modeling; (3) paper-based computational thinking tasks; (4) computational modeling of the water runoff phenomenon; and (5) engineering design, where students use their computational models developed in SPICE’s computer-based learning environment to redesign a schoolyard while considering engineering constraints [34]. The curriculum targets NGSS performance expectations for upper elementary and middle school Earth science and engineering design, emphasizing surface water movement and human impact.
@@
SPICE leverages ECD to systematically create assessments and tasks to evaluate student learning in science, computing, and engineering. For our analysis, we picked three of the seven formative assessments (F2, F3, and F5) developed to monitor and support students’ learning across the three domains in the curriculum. The locations of the assessments in the integrated curriculum are shown at the identified markers in Figure 1.
Figure 1: SPICE curricular sequence (L items in white are lessons; F items in red are formative assessments.
@@
    a Debugging Task (F3) requiring students to check the conditional form of the rules and value expressions from the Rules Task by analyzing block-based code generated by a fictional classmate and identifying errors in the computational model; and
@@
[…]
@@
R1 Completed if statement “if rainfall is less than absorption limit.” SCI, COM
@@
R4 Completed if statement “if rainfall is equal to absorption limit.” SCI, COM
@@
R7 Completed if statement “if rainfall is greater than absorption limit.” SCI, COM
@@
TABLE I: Categorical rubric used for the Rules Task. Each “R” corresponds to a different subscore for the Rules Task and is explained in the table. The Rules Task targets the science (SCI) and computing (COM) domains.
Table I enumerates the nine possible points (subscores) for the Rules Task (R1-R9). Students receive 1 point per correct conditional statement they identify (R1, R4, R7). Within each conditional statement, students receive 1 point for correctly setting the absorption value (R2, R5, R8) and 1 point for correctly setting the runoff value (R3, R6, R9). For example, the statement “if rainfall is equal to the absorption limit, then set absorption to rainfall, and set runoff to zero” would earn 3 points (R4, R5, and R6).
Figure 2: Erroneous code presented to students during the Debugging Task. Red “D” circles correspond to the individual model errors presented in Table II.
For F3, the Debugging Task, students are asked to identify and describe the five errors present in a fictional student’s code (illustrated in Figure 2). To accomplish this, students must understand the conditional logic expressions they wrote for the Rules Task and accurately translate them into block-structured programming code using “if statements” and expressions in the SPICE environment to model the amount of absorbed water and runoff. Students learn how to use the coding blocks to build their computational water runoff models before they work on the Debugging Task. The rubric for this task is shown in Table II. This paper uses the terms “model” or “computational model” to refer to the block-based code representation and “language model” or “LLM” to refer to GPT-4 (discussed in Section IV-B).
@@
TABLE II: Categorical rubric used for the Debugging Task. Each “D” corresponds to a different subscore (i.e., code error; see Figure 2) for the Debugging Task and is explained in the figure. Like the Rules Task, the Debugging Task targets the science (SCI) and computing (COM) domains.
Table II shows that students can earn up to 5 points (subscores; one for each error they identify) for the Debugging Task. “D” values refer to the individual errors the students must identify in the model. D1 refers to the “set absorption limit (inch) of the selected material” block being erroneously placed on line 6 (it should come before the first conditional statement on line 3). D2 refers to rainfall being incorrectly compared to absorption in the “less than” condition on line 7 (it should be compared to absorption limit). D3 refers to absorption incorrectly being set to the absorption limit inside the “less than” condition on line 8 (absorption should be set to rainfall). D4 refers to the “greater than” condition being improperly set on line 10 (it should not be nested inside the “less than” condition). D5 refers to the absorption and absorption limit being swapped inside the “greater than” condition on line 11 (absorption should be set to absorption limit, not the other way around).
@@
This task utilizes a new rubric structure where students are awarded a single numerical score from 0 to 4 points. Students are assessed on their ability to determine if the reported tests allow a valid comparison between the two design solutions. Since the fictional student uses different rainfall values to compare the runoff between the designs, the comparison is not “fair” because the outcome variable is not generated using the same input for both tests. The students’ explanations are evaluated at different levels using the rubric in Figure III.
@@
TABLE III: Engineering Task rubric, targeting the science (SCI), computing (COM), and engineering (ENG) domains.
Table III shows that students are awarded 0 points if they fail to recognize that the two designs cannot be compared fairly (i.e., they answer “yes” to the question posed). Students receive one point if they identify that the two tests cannot be compared fairly (i.e., they answer “no” to the question posed) but do not provide a meaningful explanation. Students receive two points if they discuss design constraints as a reason the two tests are not comparable. Three points are awarded if the students discuss the differing rainfall values as the reason why the tests cannot be compared fairly. Four points are awarded if the students mention the differing rainfall values and explain that this results in unfair runoff comparisons.
IV Methods
IV-A Chain-of-Thought Prompting + Active Learning (CoTAL)
We have developed CoTAL (illustrated in Figure 3), a generalizable method for improving automated formative assessment scoring and aligning LLM scores and explanations with the needs of teachers and students. Our approach fosters a collaborative partnership among researchers, teachers, students, and AI, integrating ECD, human-in-the-loop prompt engineering, and stakeholder feedback to (1) score and explain students’ short-answer responses to formative assessment questions in the SPICE curriculum; and (2) iteratively refine our formative assessments, rubrics, and grading prompts based on student and teacher feedback.

### v3

# CoTAL: Human-in-the-Loop Prompt Engineering for Generalizable Formative Assessment Scoring

@@ Abstract
Large language models (LLMs) have created new opportunities to assist teachers and support student learning. While researchers have explored various prompt engineering approaches in educational contexts, the degree to which these approaches generalize across domains—such as science, computing, and engineering—remains underexplored. In this paper, we introduce Chain-of-Thought Prompting + Active Learning (CoTAL), an LLM-based approach to formative assessment scoring that (1) leverages Evidence-Centered Design (ECD) to align assessments and rubrics with curriculum goals, (2) applies human-in-the-loop prompt engineering to automate response scoring, and (3) incorporates chain-of-thought (CoT) prompting and teacher and student feedback to iteratively refine questions, rubrics, and LLM prompts. Our findings demonstrate that CoTAL improves GPT-4’s scoring performance across domains, achieving gains of up to 38.9% over a non-prompt-engineered baseline (i.e., without labeled examples, chain-of-thought prompting, or iterative refinement). Teachers and students judge CoTAL to be effective at scoring and explaining responses, and their feedback produces valuable insights that enhance grading accuracy and explanation quality.
keywords:
@@
Human-in-the-Loop, Formative Assessment, Evidence-Centered Design, ECD, Automated Short Answer Scoring, Automated Grading, Prompt Engineering, LLM, LLMs, K12 STEM
1 Introduction
@@
In K-12 STEM+C (Science, Technology, Engineering, Mathematics, and Computing) classrooms, educators foster engagement by linking scientific principles to real-world phenomena, enabling students to develop exploration, inquiry, and problem-solving skills. This provides students with opportunities to develop a foundational understanding of essential STEM+C concepts and practices [1]. Unlike single-discipline curricula, STEM+C learning requires students to integrate cross-domain concepts synergistically, linking ideas from one domain (e.g., science) to another (e.g., computing). While this approach enhances learning outcomes and promotes a deeper understanding of scientific processes, it also introduces complexities that can hinder learning [2, 3], necessitating additional guidance and support that formative assessments can provide.
@@
Formative assessments employing open-ended questions help students identify key constructs, apply concepts to reasoning and problem-solving tasks, and help develop critical thinking skills [4]. They simultaneously provide teachers with granular insights into students’ STEM+C knowledge and problem-solving abilities [5], allowing them to monitor progress and adapt instruction to student needs. Unlike summative assessments and standardized tests, which are primarily evaluative and often insensitive to the complexities of students’ learning challenges [6], formative assessments promote self-reflection in students when they face difficulties and facilitate targeted feedback that helps them refine their understanding and improve performance.
@@
Evidence-based approaches for generating formative assessments and evaluation rubrics help make explicit their alignment with curricular goals and designated standards [7]. Evidence-centered design (ECD) enables a more nuanced and flexible approach by adopting cognitive science and instructional design approaches that embed evidentiary reasoning into each stage of assessment and rubric development [6].
@@
However, challenges persist in classroom environments for generating, grading, and providing timely feedback for formative assessments that align with educational standards. Designing and grading free-response formative assessments can place excessive demands on classroom teachers, who may have limited expertise in integrating STEM+C subjects, potentially affecting their ability to create assessments and rubrics that accurately evaluate students’ interdisciplinary knowledge [1]. Students may also lack mature writing skills, necessitating significant time and effort from human graders to infer the true implications of students’ answers [8]. Consequently, research is needed to develop automated scoring systems that efficiently deliver needed feedback to facilitate STEM+C learning based on teacher preferences [9].
@@
Prompt engineering with large language models (LLMs) offers a promising solution, allowing users to adapt language models to downstream tasks by incorporating approaches like in-context learning (ICL) [10], chain-of-thought (CoT) prompting [11], and active learning [12] that obviate the need for traditional training via parameter updates, thus conserving data and computational resources often unavailable in educational settings [13]. ICL enables LLMs to “learn” [^1] We use the term “learn” to indicate that the LLM uses examples to guide its predictions, not to suggest that ICL involves learning via parameter updates, which it does not. from few-shot examples in the prompt during inference. CoT extends ICL by augmenting labeled few-shot examples with step-by-step reasoning chains to provide more explicit guidance to the LLM for scoring student answers [14, 15]. Active learning is a human-in-the-loop approach to improving model training, where the human serves as an oracle, selecting additional instances to label for the next training iteration to enhance performance and robustness [12].
@@
In this paper, we leverage human-in-the-loop prompt engineering to realize a stakeholder-AI partnership [^2] For this paper, “stakeholders” refers to teachers, students, and researchers., enriching assessment and feedback within an NGSS-aligned (Next Generation Science Standards [16]) Earth science curricular unit [17, 18]. Building on our previous work on automated scoring for short-answer science assessments [19], we propose a generalizable approach: Chain-of-Thought Prompting + Active Learning (CoTAL) that employs LLMs to automate scoring and feedback for formative assessments across multiple domains. We define generalizability as the ability to grade formative short-answer responses varying by question type (e.g., concept definitions, process descriptions, comparisons, explanations), rubric structure (e.g., multi-label vs. multi-class), and content domain (science, computing, engineering).
To evaluate CoTAL, we address the following research questions:
@@
2 Background
2.1 Evidence-Centered Design
ECD structures assessments around evidentiary reasoning. Domain, Evidence, and Task models specify the knowledge, skills, and abilities (KSAs) aligned with standards (e.g., NGSS [16]) and expertise; identify observable indicators of mastery; and design activities that elicit the required evidence [6]. This alignment supports formative assessments and rubrics that faithfully track student understanding against curricular goals.
LLMs enable evidence elicitation during automated formative assessment grading via CoT prompting. Beyond assigning scores, LLMs can quote student responses, tie them to assessment questions and rubric items, and pinpoint areas of proficiency and difficulty. This alignment grounds LLM reasoning, minimizes hallucinations, and supports accurate scoring and mastery evidence generation.
While substantial work has explored ECD-driven formative assessment [23, 24, 6], few studies have applied ECD principles to automating formative assessment scoring and feedback. Hao (2025) [25] takes a critical first step by examining the challenges and opportunities of integrating ECD into LLM systems, advocating for “continuous research, critical evaluation, and open collaboration to ensure responsible, effective integration” of LLMs into assessment practices. However, this work stops short of proposing a concrete implementation. Zapata et al. (2024) [26] investigate the integration of LLMs and ECD for conversation-based assessment scoring, explicitly calling for human-in-the-loop approaches to mitigate fairness and bias issues in LLMs. Their study focused primarily on the prompt engineering process and did not evaluate ECD-driven LLM feedback with classroom learners.
2.2 Automated Formative Assessment Scoring and Feedback
@@
Previously, automated assessment scoring methods have incorporated data augmentation [27, 28, 29], next sentence prediction [30], domain adaptation and supervised fine-tuning [31], prototypical neural networks [32], cross-prompt fine-tuning [33], and reinforcement learning [34] to improve grading accuracy. However, these methods have largely focused on more structured grading tasks in mathematics and computer science [35, 36], where open-ended responses are less prevalent than in science domains [19]. These approaches have achieved varying degrees of success but often fail to provide comprehensive insights into their scoring decisions.
More recently, researchers have explored the use of LLMs for automated scoring in science [37]. Golchin et al. (2024) [14] adapted zero-shot chain-of-thought prompting to grade writing assignments in Coursera courses on Astronomy and Astrobiology, benchmarking GPT-4 against instructor and peer evaluations. Impey et al. (2025) [15] assessed GPT-4’s rubric-based grading reliability on similar MOOC assignments, finding its scores closely aligned with both instructors and peers. While Impey et al. (2025) caution that rubric interpretation can vary between humans and GPT-4 and note that findings from low-stakes MOOC environments may not generalize to high-stakes classroom settings, Golchin et al. (2024) report degraded performance on tasks requiring creative or speculative reasoning—underscoring the need for human-in-the-loop evaluations conducted in situ. Pinto et al. (2023) [38] used ChatGPT to grade open-ended questions in technical training, experimenting with different levels of determinism (i.e., temperature settings). However, their study was conducted strictly in a zero-shot setting and focused on adult technical training rather than K-12 STEM learners.
Researchers have also begun to leverage the rich feedback capabilities offered by LLMs. Matelsky et al. (2023) [39] present a domain-agnostic, open-source widget that uses LLMs to automate personalized feedback on students’ open-ended responses. However, without human validation, the tool’s effectiveness hinges on the quality and completeness of the criteria provided by the user (i.e., the instructor), and it has not yet been empirically validated in real-world classrooms. Meyer et al. (2024) [40] used LLMs to bring evidence-based feedback into the classroom, investigating whether AI-generated feedback improves secondary students’ essay revisions, motivation, and emotions compared to no feedback. However, their study focuses on English-as-a-foreign-language students rather than STEM learners.
2.3 Human-in-the-Loop Prompt Engineering
Unlike algorithmic prompt engineering, human-in-the-loop approaches [19, 41] integrate an important human perspective into the design and optimization pipeline that enables users to influence LLM outputs rather than relying solely on algorithmic decisions [42, 41]. This creates a human-AI collaboration [43, 44] that enhances LLM alignment with user preferences and ensures the contextual relevance of the generated output by incorporating nuanced domain knowledge that algorithmic methods may overlook. Furthermore, this approach helps mitigate LLM errors and hallucinations by leveraging human expertise to validate and refine generated responses, fostering more accurate, ethical, and trustworthy LLM interactions.
Beyond Cohn et al. [19, 41], Lee et al. (2024) [45] showed that CoT prompting with contextual item stems and rubrics improves LLM scoring of middle school science assessments by emphasizing domain-specific reasoning. Their human-in-the-loop procedure, WRVRT (Writing, Reviewing, Validating, Revising, Testing), combines CoT prompting with iterative refinement to enhance accuracy and explainability. However, their work is limited to science, leaving cross-subject generalizability untested; moreover, it centers on prompt engineering without incorporating ECD principles or integrating teacher and student feedback to refine assessments, rubrics, or prompts.
2.4 Active Learning
The term active learning is formally defined in both educational and AI contexts. In education, active learning is defined by Bonwell and Eison (1991) [46] as “instructional activities involving students in doing things and thinking about what they are doing,” emphasizing a shift from passive information reception to skill development through meaningful engagement (e.g., group work and projects) that fosters discussion and reflection.
In AI, Cohn et al. (1994) [12] define active learning as “any form of learning [i.e., training] in which the learning program has some control over the inputs on which it trains.” Rather than passively consuming labeled examples, active learning algorithms selectively query a human “oracle” for labels on the most informative instances to optimize the training process. This approach ensures models are trained on impactful data, learning from fewer examples and inducing more precise decision boundaries.
In our work, we adopt the AI definition of active learning, establishing it as a human-in-the-loop prompt engineering approach to iteratively refine prompts and augment them with few-shot examples to improve in-context learning by: (1) testing prompts on a validation set; (2) identifying trends in scoring deficiencies (i.e., factors causing LLM misalignment with human judgment); and (3) incorporating targeted few-shot examples using chain-of-thought (CoT) prompting to address those deficiencies and correct misalignment, following the approach introduced by Cohn et al. (2024) [19].
Other researchers have proposed a related strategy. i.e., active prompting [47, 48], which leverages uncertainty metrics (e.g., entropy or margin) to identify examples with the least confident predictions. The top-$k$ most ambiguous cases are sent to human annotators for labeling or corrective feedback, which are then incorporated directly into the prompt as demonstrative examples. However, unlike active learning, active prompting emphasizes aggregate uncertainty rather than targeting specific misalignment patterns responsible for LLM disagreement. In contrast, our work addresses specific LLM misalignments in cases where discernible trends emerge during validation set evaluation, as elaborated in Section 4.1.
2.5 Research Gaps
While researchers have employed a variety of prompt engineering techniques to improve LLM performance on formative assessment scoring [49], these approaches typically focus on a single domain and dataset, leaving their generalizability largely unevaluated. Moreover, they often neglect stakeholder input in the development and refinement of their methods. A notable gap remains in automated formative assessment scoring approaches that (1) integrate ECD principles with prompt engineering, (2) provide transparent explanations of scoring decisions to teachers and students, (3) demonstrate effectiveness across multiple domains, and (4) incorporate stakeholder feedback for iterative refinement.
@@ 2.5 Research Gaps
In this paper, we address these gaps by proposing a generalizable, stakeholder-informed framework that combines ECD with human-in-the-loop prompt engineering (i.e., active learning) to enable transparent, accurate, and adaptable formative assessment scoring. Our work builds on prior research by explicitly grounding our methodology in assessment design theory, engaging both teachers and students in the refinement process, and evaluating performance across science, engineering, and computing domains in real-world classroom settings.
3 SPICE Curriculum
@@
SPICE [^3] https://spiceprojects.org/ (Science Projects Integrating Computing and Engineering) is a three-week middle school curriculum unit in which students redesign a schoolyard using surface materials that minimize post-rainstorm runoff while satisfying design constraints (e.g., cost, accessibility). The problem-based curriculum comprises five units: (1) physical experiments; (2) conceptual modeling; (3) paper-based computational thinking tasks; (4) computational modeling of water runoff; and (5) engineering design using the computational model to evaluate solutions. Students build computational models in SPICE’s computer-based environment and apply them to redesign the schoolyard under engineering constraints [50]. The curriculum targets NGSS performance expectations for upper elementary and middle school Earth science and engineering design, emphasizing surface water movement and human impact.
@@
SPICE leverages ECD to systematically create assessments and tasks to evaluate student learning in science, computing, and engineering. For our analysis, we picked three of the seven formative assessments: (1) F2(conceptual modeling in science); (2) F3(computational modeling of science phenomena); and (3) F5(testing engineering designs), developed to monitor and support students’ learning across the three domains in the curriculum. The locations of the assessments in the integrated curriculum are shown at the identified markers in Figure 1.
Figure 1: SPICE curricular sequence (L items in white are lessons; F items in red are formative assessments. “CT” refers to computational thinking, which we use synonymously with “computing” for the purposes of this paper.
@@
    a Debugging Task (F3) requiring students to check the conditional form of the rules and value expressions from the Rules Task by analyzing block-based code generated by a fictional classmate to generate a computational science model of the runoff phenomena and identifying errors in the computational model; and
@@
The Rules Task nudges students to translate their learned intuitions about the conservation of matter principle into a quantitative relation between variables (total rainfall, total absorption, absorption limit, and total runoff). Students express this relation by considering three scenarios, i.e., when rainfall is greater than, less than, and equal to the surface absorption limit. Students are then asked to express these three scenarios as conditional logic expressions defining absorption and runoff values. For example, if rainfall<<absorption limit of the ground material, then absorption = rainfall and runoff = 0. Later, students must recall these conditional logic expressions to construct their computational models. We identify the Rules Task rubric as categorical, as there is a specific structure to the response, and students receive 1 point for including each required component per rule. The rubric for the Rules Task appears in Table 1.
@@
R1 Completed if statement “if rainfall is << absorption limit.” SCI, COM
@@
R4 Completed if statement “if rainfall == to absorption limit.” SCI, COM
@@
R7 Completed if statement “if rainfall >> than absorption limit.” SCI, COM
@@
Table 1: Categorical rubric used for the Rules Task. Each “R” corresponds to a different subscore for the Rules Task and is explained in the table. The Rules Task targets the science (SCI) and computing (COM) domains (with an emphasis on science).
Table 1 enumerates the nine possible points (subscores) for the Rules Task (R1-R9). Students receive 1 point per correct conditional statement they identify (R1, R4, R7). Within each conditional statement, students receive 1 point for correctly setting the absorption value (R2, R5, R8) and 1 point for correctly setting the runoff value (R3, R6, R9). For example, the statement “if rainfall is equal to the absorption limit, then set absorption to rainfall, and set runoff to zero” would earn 3 points (R4, R5, and R6).
Figure 2: Erroneous code presented to students during the Debugging Task. Red “D” circles correspond to the individual model errors presented in Table 2.
For F3, the Debugging Task, students are asked to identify and describe the five errors present in a fictional student’s code (illustrated in Figure 2). To accomplish this, students must understand the conditional logic expressions they wrote for the Rules Task and accurately translate them into block-structured programming code using “if statements” and expressions in the SPICE environment to model the amount of absorbed water and runoff. Students learn how to use the coding blocks to build their computational water runoff models before they work on the Debugging Task. The rubric for this task is shown in Table 2.
@@
Table 2: Categorical rubric used for the Debugging Task. Each “D” corresponds to a different subscore (i.e., code error; see Figure 2) for the Debugging Task and is explained in the figure. Like the Rules Task, the Debugging Task targets the science (SCI) and computing (COM) domains (but with an emphasis on computing).
Table 2 shows that students can earn up to 5 points (subscores; one for each error they identify) for the Debugging Task. “D” values refer to the individual errors the students must identify in the model. D1 refers to the “set absorption limit (inch) of the selected material” block being erroneously placed on line 6 (it should come before the first conditional statement on line 3). D2 refers to rainfall being incorrectly compared to absorption in the “less than” condition on line 7 (it should be compared to the absorption limit). D3 refers to absorption incorrectly being set to the absorption limit inside the “less than” condition on line 8 (absorption should be set to rainfall). D4 refers to the “greater than” condition being improperly set on line 10 (it should not be nested inside the “less than” condition). D5 refers to the absorption and absorption limit being swapped inside the “greater than” condition on line 11 (absorption should be set to absorption limit, not the other way around).
@@
This task utilizes a new rubric structure where students are awarded a single numerical score from 0 to 4 points. Students are assessed on their ability to determine if the reported tests allow a valid comparison between the two design solutions. Since the fictional student uses different rainfall values (i.e., inputs) to compare the runoff between the designs, the comparison is not “fair” because the outcome variable is not generated using the same input for both tests. The students’ explanations are evaluated at different levels using the rubric in Figure 3.
@@
Table 3: Engineering Task rubric, targeting the science (SCI), computing (COM), and engineering (ENG) domains (with an emphasis on engineering).
Table 3 shows that students are awarded 0 points if they fail to recognize that the two designs cannot be compared fairly (i.e., they answer “yes” to the question posed). Students receive one point if they identify that the two tests cannot be compared fairly (i.e., they answer “no” to the question posed) but do not provide a meaningful explanation. Students receive two points if they discuss design constraints as a reason that the two tests are not comparable. Three points are awarded if the students discuss the differing rainfall values as the reason why the tests cannot be compared fairly. Four points are awarded if the students mention the differing rainfall values and explain that this results in unfair runoff comparisons.
4 Methods
4.1 Chain-of-Thought Prompting + Active Learning (CoTAL)
We have developed CoTAL (illustrated in Figure 3), as a generalizable method for improving automated formative assessment scoring and aligning LLM scores and explanations with the needs of teachers and students. Our approach fosters a collaborative partnership among researchers, teachers, students, and AI, integrating ECD, human-in-the-loop prompt engineering, and stakeholder feedback to (1) score and explain students’ short-answer responses to formative assessment questions in the SPICE curriculum; and (2) iteratively refine our formative assessments, rubrics, and grading prompts based on student and teacher feedback.